The kinetic energy of an object of mass 2kg is 16J. What is the linear momentum?
Answer #1

mass m = 2 kg

kinetic energy = E = 16 J

MOMENTUM P = ?

           E = P^2/ 2M =

P= sqt( 2m E )= 8 Kg m/sec